# 1NLPR, MAIS, CASIA 2UCAS 3PKU 4WHU 5ByteDance

### v2

@@ Abstract
While Multimodal Large Language Models (MLLMs) excel at holistic understanding, they struggle in capturing the dense world with complex scenes, requiring fine-grained analysis of intricate details and object inter-relationships. Region-level MLLMs have been a promising step. However, previous attempts are generally optimized to understand given regions in isolation, neglecting crucial global contexts. To address this, we introduce Grasp Any Region (GAR) for comprehensive region-level visual understanding. Empowered by an effective RoI-aligned feature replay technique, GAR supports (1) precise perception by leveraging necessary global contexts, and (2) modeling interactions between multiple prompts. Together, it then naturally achieves (3) advanced compositional reasoning to answer specific free-form questions about any region, shifting the paradigm from passive description to active dialogue. Moreover, we construct GAR-Bench, which not only provides a more accurate evaluation of single-region comprehension, but also, more importantly, measures interactions and complex reasoning across multiple regions. Extensive experiments have demonstrated that GAR-1B not only maintains the state-of-the-art captioning capabilities, e.g., outperforming DAM-3B +4.5 on DLC-Bench, but also excels at modeling relationships between multiple prompts with advanced comprehension capabilities, even surpassing InternVL3-78B on GAR-Bench-VQA. More importantly, our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B on VideoRefer-Bench, indicating its strong capabilities can be easily transferred to videos.
@@ 1 Introduction
The ambition of Multimodal Large Language Models (MLLMs) is to endow machines with human-like abilities to perceive, interpret, and reason about the dense visual world [60, 22, 21]. To date, renowned state-of-the-art models [1, 52, 50, 10, 31, 32, 33] have made remarkable strides, excelling in answering general questions about an entire image. However, this global-level perception struggles with the dense understanding of cluttered environments, intricate object details, and the complex interplay between multiple entities.
@@
To address the limitation of global perception, several previous works [3, 60, 64, 22, 25] argue for a paradigm shift to region-level MLLMs. Specifically, they equip MLLMs with promptable and fine-grained interactions to achieve targeted region-level understanding, using boxes [64, 3] or masks [60, 22]. This mechanism transforms the model from a passive observer of the entire scene into an active participant capable of deep, localized analysis. Conventional region MLLMs [22, 25] mainly focus on the ability to generate a descriptive caption for a single region, and thus model architectures are generally optimized to understand a given region in isolation. This design often neglects crucial global context, e.g., misidentifying a frog-shaped slipper as a real frog in Figure 2a. Alternatives [60, 58] that employ pooled local features suffer from insufficient details. Therefore, a unified framework that can simultaneously resolve these issues to facilitate more sophisticated and interactive capabilities remains a significant area for investigation. To this end, we propose Grasp Any Region (GAR) for comprehensive region understanding. As shown in Figure 2, key features include:
@@
(1) Precise Perception. Thanks to the leverage of necessary global contexts, GAR achieves a more precise perception of given regions, which is the fundamental capability for region MLLMs. As shown in Figure 2a by aggregating information from the broader, unmasked scene, our GAR manages to generate much more accurate descriptions than previous crop-based approaches [22].
@@
(2) Interactions between Multiple Prompts. Our framework moves beyond the prevailing single-prompt paradigm, which treats every region of interest as an isolated entity. As illustrated in Figures 2b and 2c, GAR manages to model relationships between an arbitrary number of prompts.
@@
(3) Advanced Compositional Reasoning Capabilities. Empowered with the aforementioned features, GAR is naturally equipped with advanced compositional reasoning capabilities, allowing it to answer any specific free-form questions, e.g., recognizing non-entities shown in Figure 2c.
@@
To achieve these capabilities, effectively encoding global contexts becomes equally crucial as local detailed features. To this end, we propose an RoI-aligned feature replay technique. Specifically, GAR first encodes the full, uncropped image (together with the mask prompt) with AnyRes [27]. Subsequently, RoI-Align [16] is employed to gather relevant features directly from the global feature map. Those gathered features are inherently context-aware, providing sufficient local details while maintaining global information simultaneously. Please refer to Figure 3 for the detailed pipeline.
@@
Empirically, shown in Figure 1, our GAR-1B not only outperforms DAM-3B [22] and PAM-3B [25] on detailed captioning benchmarks [22, 58, 24], but also excels in general multimodal benchmarks [51, 43, 53, 4]. Interestingly, it even outperforms large-scale models like InternVL3-78B [66] on GAR-Bench, demonstrating its advanced comprehension capability in modeling interactions between multiple prompts. More importantly, our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B on VideoRefer-Bench, indicating its strong comprehension capabilities can be easily transferred to videos. We hope our work inspires the community to develop MLLMs that can perceive and understand the dense visual world more effectively.
@@ 2 Related Works
Multimodal Large Language Models (MLLMs). Typical MLLMs [26, 20, 27, 1, 66, 52, 19, 46, 47, 42, 56, 57, 19] project visual features extracted from pre-trained visual encoders [35, 62] to LLM for understanding multimodal contents. However, these models usually lack precise localization capabilities [22, 25] and struggle to understand specific regions. One potential solution is to “think with images” [33, 48, 45]. But these agentic models require complex multi-turn conversations, while we mainly focus on precise perception within a single-turn dialogue.
Region-Level MLLMs. Different from conventional image-level comprehension, localized understanding requires MLLMs to capture regional attributes. Previous methods either utilize visual markers [54], bounding boxes [64, 3, 58, 38], or segmentation masks [60, 22], to represent regions-of-interests within an image. We simply regard masks as visual prompts, since masks have less ambiguity than other representations. However, previous approaches only support a single visual prompt, and often neglect global context. GAR is designed for modeling the relationship between an arbitrary number of visual prompts while effectively maintaining crucial global context.
Benchmarks for Region-Level Understanding. Typical region-level benchmarks only evaluate the caption quality for single prompt using conventional language-based captioning metrics [58, 60, 65, 14, 38], model-based similarities [2, 60], and LLM-Judged accuracies without the need for reference captions [22]. GAR-Bench is to systematically evaluate the comprehension capabilities with multiple visual prompts. It contains a caption protocol to measure the correctness of descriptions for the relation between visual prompts, and a VQA protocol to evaluate both the basic understanding capability for specific regions, e.g., color and shape, and advanced compositional reasoning abilities for multiple regions.
@@ 3 Grasp Any Region
We start from the task formulation in § 3.1. Subsequently, we introduce our model architecture and training data pipeline in § 3.2 and § 3.3, respectively. Finally, we introduce our benchmark designs in § 3.4 to systematically evaluate region-level comprehension capabilities.
@@ 3.2 Model Architecture
Prompt Encoding and Integration. To integrate spatial guidance into the vision backbone, we introduce a lightweight prompt encoding mechanism similar to [22] and [40]. The input binary mask, which specifies the region(s) of interest, is first processed by a simple convolutional block [18] to produce a mask embedding. This zero-initialized [63] mask embedding is then added to ViT’s [12] patch embeddings.
RoI-aligned Feature Replay. To simultaneously provide sufficient local details and maintain necessary global context, we introduce the RoI-aligned feature replay technique. Specifically, our model processes the full, uncropped image (with the encoded mask prompt) with AnyRes [27], producing a global feature map that is rich in contextual information. Based on the input mask, we then derive a corresponding bounding box for the region of interest and employ RoI-Align [16] to gather the relevant feature vectors directly from the global feature map. Because the features are extracted from a feature map that was computed over the entire image, they are inherently context-aware, which elegantly avoids the pitfalls of local-only processing in [22]. At the same time, it provides the subsequent language model with a sufficiently detailed, high-resolution representation of the prompted region, enabling it to perform fine-grained understanding. This replay of context-rich features allows GAR to simultaneously “zoom in” on detail without “losing sight” of the bigger picture. Ablations of this design can be found in Table 8, where we demonstrate that this design is capable of both (1) providing sufficient local details and (2) preserving global contexts.
@@ 3.3 Training Data Pipeline
To enhance model capabilities from basic object recognition with single region to complex relational reasoning with multiple regions, we design a multi-stage process to generate a large-scale, high-quality dataset, as illustrated in Figure 4. Ablations of each round can be found in Table 11. Prompts for each stage can be found in § G.
Round 1: Enhance Recognition Capability. Initially, we start from the Describe Anything-1.5M dataset [22]. However, we observe deficiencies in its fine-grained recognition capability, limiting the quality of generated captions for more complex scenarios. To address this, we integrated images and masks provided by [40], which is a subset of ImageNet-21K [11], an extremely fine-grained classification dataset and renowned for its detailed and extensive category labels. We employ the seed captioner to generate descriptions and then utilize an LLM to validate these generated captions against the ground-truth categories, resulting in a refined fine-grained dataset of 456K samples. We utilize both datasets to train a fine-grained captioner.
Round 2: Supporting Multiple Prompts. To further enable understanding multiple prompts, we incorporated the Panoptic Scene Graph (PSG) dataset [55], which is rich in relational information. We first query the fine-grained captioner to generate a detailed description for each region. Subsequently, we regard Qwen2.5-72B [41] as the LLM-Merger, together with the original annotations provided by the PSG dataset [55], to generate: (1) 144K rich object descriptions that explicitly integrate relational context, (2) 144K question-answering pairs designed to probe the understanding of complex relationships, and (3) 126K multiple-choice questions. We construct a relation dataset with 414K samples in total during this stage.
@@ 4 Experiments
Owing to page limitations, we only present the key properties in this section. For implementation details, comparative baselines, and ablation studies, please refer to § C.
@@
| Qwen2.5-VL-3B [1] | 34.4 | 29.0 | 25.0 | 34.5 | 30.6 | 43.8 | 26.2 | 44.6 | ||||||||||||
| Qwen2.5-VL-7B [1] | 41.7 | 39.1 | 40.6 | 44.8 | 27.8 | 59.4 | 36.1 | 40.6 | ||||||||||||
| Qwen2.5-VL-32B [1] | 50.9 | 46.4 | 53.1 | 41.4 | 30.6 | 71.9 | 36.1 | 58.4 | ||||||||||||
| Qwen2.5-VL-72B [1] | 52.8 | 46.4 | 50.0 | 65.5 | 33.3 | 68.8 | 44.3 | 57.4 | ||||||||||||
| InternVL3-2B [66] | 35.1 | 30.4 | 21.9 | 48.3 | 38.9 | 48.4 | 26.2 | 38.6 | ||||||||||||
| InternVL3-8B [66] | 38.9 | 36.2 | 37.5 | 58.6 | 41.7 | 51.6 | 27.9 | 33.6 | ||||||||||||
| InternVL3-38B [66] | 46.5 | 39.1 | 40.6 | 51.7 | 55.6 | 60.9 | 36.1 | 47.5 | ||||||||||||
| InternVL3-78B [66] | 50.5 | 44.9 | 54.7 | 58.6 | 61.1 | 53.1 | 47.5 | 45.5 | ||||||||||||
@@
| Qwen2.5-VL-3B [1] | 22.5 | 9.3 | 34.6 | ||||||
| Qwen2.5-VL-7B [1] | 32.4 | 12.4 | 50.5 | ||||||
| Qwen2.5-VL-32B [1] | 36.8 | 17.5 | 54.3 | ||||||
| InternVL3-2B [66] | 29.4 | 14.4 | 43.0 | ||||||
| InternVL3-8B [66] | 33.8 | 11.3 | 54.2 | ||||||
| InternVL3-38B [66] | 45.1 | 29.9 | 58.9 | ||||||
@@
| GPT4RoI-7B [64] | 26.3 | 6.5 | 46.2 |
| Shikra-7B [3] | 22.2 | 2.7 | 41.8 |
| Ferret-7B [58] | 22.4 | 6.4 | 38.4 |
| RegionGPT-7B [14] | 27.2 | 13.0 | 41.4 |
| VP-SPHINX-13B [24] | 22.5 | 11.7 | 33.2 |
| DAM-3B [22] | 64.5‡ | 47.2‡ | 81.8‡ |
@@
Advanced comprehension requires precisely modeling complex relationships between multiple prompts. To evaluate this capability, we conducted a comprehensive comparison on our GAR-Bench-VQA. As demonstrated in Table 1, GAR-8B achieves an impressive overall score of 54.5, surpassing even the powerful, private, state-of-the-art non-thinking model, GPT-4o [31]. Furthermore, the efficiency and effectiveness of our approach are highlighted by GAR-1B. Despite its significantly smaller size, it scores 50.6 overall, outperforming large-scale public models like InternVL3-78B [66]. This advantage is particularly evident in fine-grained perception tasks, where GAR-1B and GAR-8B achieve “Texture” scores of 69.0 and 75.9, respectively.
@@
Detailed localized captioning requires generating detailed descriptions for given regions with multiple sentences. We benchmark our GAR models on a series of challenging datasets, and the results consistently demonstrate their state-of-the-art capabilities. As shown in Table 3, on our GAR-Bench-Cap, GAR-1B and GAR-8B achieve the highest overall scores of 57.5 and 62.2, respectively, even exceeding that of powerful private models like Gemini-2.5-Pro [10]. This superiority is further confirmed on the DLC-Bench [22] in Table 3, where GAR-1B and GAR-8B again outperform top models like DAM-3B using either LLaMA3.1 [13] or GPT-4o [31] as the judge. The zero-shot performance of our models on Ferret-Bench [58] and MDVP-Bench [24], detailed in Table 5, is particularly noteworthy. On both benchmarks, our GAR emerges as the top-performing model across every single category. Specifically on MDVP-Bench, our models show a commanding lead, with GAR-8B achieving a score of 178.6 on natural images, a result that is substantially higher than any competitor. Collectively, these comprehensive evaluations across multiple benchmarks unequivocally establish GAR as the new state-of-the-art for producing rich, accurate, and detailed localized captions.
Open-class category-level image recognition requires the model to recognize the category of the object and part entities. We evaluate this capability in Table 5. Our GAR-8B demonstrates a significant leap in performance, establishing a new state-of-the-art. It consistently outperforms all prior methods across every metric, achieving top scores of 93.6 semantic similarity and 88.7 semantic IoU on LVIS [15], and 95.5 semantic similarity and 91.8 semantic IoU on PACO [36]. This indicates its superior ability in both semantic understanding and precise localization. These results demonstrate the effectiveness of GAR for complex recognition tasks, showcasing its robust performance in identifying a diverse range of object categories.
Extension to videos is straightforward. Similar to [22], we simply extend our GAR models to videos and evaluate them on VideoRefer-Bench [61] and VideoRefer-Bench [61] in Table 6 and Table 7, respectively. We uniformly sample 16 frames to represent a video. Our GAR-8B surpasses DAM-8B [22] under the zero-shot setting. More importantly, as demonstrated in Table 7, our our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B, demonstrating its strong comprehension capabilities can be easily transferred to videos. However, as our models are actually trained with images, they get reasonably low scores on temporally related tasks, e.g., temporal description (TD) in Table 6 and future predictions in Table 7.
@@
| LLaVA-OneVison-7B [20] | 2.12 | 2.62 | 1.58 | 2.19 | 2.07 | 2.48 | 3.09 | 1.94 | 2.50 | 2.41 |
| Qwen2-VL-7B [49] | 2.39 | 2.97 | 2.24 | 2.03 | 2.31 | 2.55 | 3.30 | 2.54 | 2.22 | 2.12 |
| InternVL2-26B [6] | 2.84 | 3.55 | 2.99 | 2.57 | 2.25 | 3.20 | 4.08 | 3.35 | 3.08 | 2.28 |
@@
| Elysium-7B [44] | 1.57 | 2.35 | 0.30 | 0.02 | 3.59 | – | – | – | – | – |
| Ferret-7B [58] | 2.18 | 3.08 | 2.01 | 1.54 | 2.14 | 2.23 | 3.20 | 2.38 | 1.97 | 1.38 |
| Osprey-7B [60] | 2.34 | 3.19 | 2.16 | 1.54 | 2.45 | 2.41 | 3.30 | 2.66 | 2.10 | 1.58 |
| Artemis-7B [34] | – | – | – | – | – | 2.26 | 3.42 | 1.34 | 1.39 | 2.90 |
| DAM-8B [22] | – | – | – | – | – | 3.34 | 4.45 | 3.30 | 3.03 | 2.58 |
@@
| InternVL2-26B [6] | 65.0 | 58.5 | 63.5 | 53.4 | 88.0 | 78.9 | |||||||||||||||||
| Qwen2-VL-7B [49] | 66.0 | 62.0 | 69.6 | 54.9 | 87.3 | 74.6 | |||||||||||||||||
| LLaVA-OneVision-7B [20] | 67.4 | 58.7 | 62.9 | 64.7 | 87.4 | 76.3 | |||||||||||||||||
| GPT-4o [31] | 71.3 | 62.3 | 74.5 | 66.0 | 88.0 | 73.7 | |||||||||||||||||

### v3

# Grasp Any Region: Towards Precise, Contextual Pixel Understanding for Multimodal LLMs

@@ Abstract
While Multimodal Large Language Models (MLLMs) excel at holistic understanding, they struggle in capturing the dense world with complex scenes, requiring fine-grained analysis of intricate details and object inter-relationships. Region-level MLLMs have been a promising step. However, previous attempts are generally optimized to understand given regions in isolation, neglecting crucial global contexts. To address this, we introduce Grasp Any Region (GAR) for comprehensive region-level visual understanding. Empowered by an effective RoI-aligned feature replay technique, GAR supports (1) precise perception by leveraging necessary global contexts, and (2) modeling interactions between multiple prompts. Together, it then naturally achieves (3) advanced compositional reasoning to answer specific free-form questions about any region, shifting the paradigm from passive description to active dialogue. Moreover, we construct GAR-Bench, which not only provides a more accurate evaluation of single-region comprehension, but also, more importantly, measures interactions and complex reasoning across multiple regions. Extensive experiments have demonstrated that GAR-1B not only maintains the state-of-the-art captioning capabilities, e.g., outperforming DAM-3B +4.5 on DLC-Bench, but also excels at modeling relationships between multiple prompts with advanced comprehension capabilities, even surpassing InternVL3-78B on GAR-Bench-VQA. More importantly, our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B on VideoRefer-Bench, indicating its strong capabilities can be easily transferred to videos. The code is available at https://github.com/Haochen-Wang409/Grasp-Any-Region.
@@ 1 Introduction
The ambition of Multimodal Large Language Models (MLLMs) is to endow machines with human-like abilities to perceive, interpret, and reason about the dense visual world (Yuan et al., 2024; Lian et al., 2025; Li et al., 2025). To date, renowned state-of-the-art models (Bai et al., 2025; Wu et al., 2024; Wang et al., 2025e; DeepMind, 2025b; OpenAI, 2024a; b; 2025) have made remarkable strides, excel in generating holistic descriptions and answering general questions about an entire image. However, this global-level perception struggles with the dense understanding of cluttered environments, intricate object details, and the complex interplay between multiple entities.
@@
To address the limitation of global perception, several previous works (Chen et al., 2023; Yuan et al., 2024; Zhang et al., 2024a; Lian et al., 2025; Lin et al., 2025b) argue for a paradigm shift to region-level MLLMs. Specifically, they equip MLLMs with promptable and fine-grained interactions to achieve targeted region-level understanding, using boxes (Zhang et al., 2024a; Chen et al., 2023) or masks (Yuan et al., 2024; Lian et al., 2025). This mechanism transforms the model from a passive observer of the entire scene into an active participant capable of deep, localized analysis. However, effectively balancing global scene context with fine-grained local details remains challenging, which serves as a fundamental trade-off in region-level MLLMs. Conventional methods (Yuan et al., 2024; You et al., 2023) that employ pooled local features suffer from insufficient details, while recent models (Lian et al., 2025; Lin et al., 2025b) mainly focus on the ability to generate a descriptive caption for a single region, and thus model architectures are generally optimized to understand a given region in isolation. This design often neglects crucial global context, e.g., misidentifying a frog-shaped slipper as a real frog in Figure 1a. To this end, we propose Grasp Any Region (GAR) for comprehensive and detailed region understanding. As shown in Figure 1, key features include:
@@
(1) Precise Perception. Thanks to the leverage of necessary global contexts, GAR achieves a more precise perception of given regions, which is the fundamental capability for region MLLMs. As shown in Figure 1a by aggregating information from the broader, unmasked scene, our GAR manages to generate much more accurate descriptions than previous crop-based approaches (Lian et al., 2025).
@@
(2) Interactions between Multiple Prompts. GAR moves beyond the prevailing single-prompt paradigm (Lian et al., 2025), which treats every region of interest as an isolated entity. As illustrated in Figures 1b and 1c, GAR manages to model relationships between an arbitrary number of prompts.
@@
(3) Advanced Compositional Reasoning Capabilities. Empowered with the aforementioned features, GAR is naturally equipped with advanced compositional reasoning capabilities, allowing it to answer any specific free-form questions.
@@
To achieve these capabilities, effectively encoding global contexts becomes equally crucial as local detailed features. To this end, we propose an RoI-aligned feature replay technique. Specifically, GAR first encodes the full, uncropped image (together with the mask prompt) with AnyRes (Liu et al., 2024). Subsequently, RoI-Align (He et al., 2017) is employed to gather relevant features directly from the global feature map. Those gathered features are inherently context-aware, providing sufficient local details while maintaining global information simultaneously. Please refer to Figure 3 for the detailed pipeline.
@@
Empirically, shown in Figure 2, our GAR-1B not only outperforms DAM-3B (Lian et al., 2025) and PAM-3B (Lin et al., 2025b) on detailed captioning benchmarks (Lian et al., 2025; You et al., 2023; Lin et al., 2025a), but also excels in general multimodal benchmarks (Wu & Xie, 2024; Tong et al., 2024b; xAI, 2024; Chen et al., 2024a). Interestingly, it even outperforms large-scale models like InternVL3-78B (Zhu et al., 2025) on GAR-Bench, demonstrating its advanced comprehension capability in modeling interactions between multiple prompts. More importantly, our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B on VideoRefer-Bench, indicating its strong comprehension capabilities can be easily transferred to videos. We hope our work inspires the community to develop MLLMs that can perceive and understand the dense visual world more effectively.
@@ 2 Related Works
Multimodal Large Language Models (MLLMs). Typical MLLMs (Liu et al., 2023; Li et al., 2024; Liu et al., 2024; Bai et al., 2025; Zhu et al., 2025; Wu et al., 2024; Lei et al., 2025; Wang et al., 2025b; c; Tong et al., 2024a; Yang et al., 2024a; b; 2025b; 2025a; 2026) project visual features extracted from pre-trained visual encoders (Radford et al., 2021; Zhai et al., 2023) to LLM for understanding multimodal contents. However, these models usually lack precise localization capabilities (Lian et al., 2025; Lin et al., 2025b) and struggle to understand specific regions. One potential solution is to “think with images” (OpenAI, 2025; Wang et al., 2025d; a). But these agentic models require complex multi-turn conversations, while we mainly focus on precise perception within a single-turn dialogue.
Region-Level MLLMs. Different from conventional image-level comprehension, localized understanding requires MLLMs to capture regional attributes. Previous methods either utilize visual markers (Yang et al., 2023), bounding boxes (Zhang et al., 2024a; Chen et al., 2023; You et al., 2023; Rasheed et al., 2024; Lee et al., 2024; Ma et al., 2024), or segmentation masks (Yuan et al., 2024; Lian et al., 2025), to represent regions-of-interests within an image. We simply regard masks as visual prompts, since masks have less ambiguity than other representations. Beyond prompt representations, effectively balancing global scene context with local details remains an open problem. Existing methods struggle to master both. Methods like DAM (Lian et al., 2025) excel at local perception but lack a holistic view of the global context. Conversely, earlier works like GPT4RoI (Zhang et al., 2024a) and GLaMM (Rasheed et al., 2024), while incorporating the global image, tend to lose crucial local details by pooling region features into single vectors. GAR is designed specifically to solve this dilemma. Based on this, GAR excels in modeling the relationship between an arbitrary number of visual prompts while effectively maintaining crucial global context and sufficient local details.
Benchmarks for Region-Level Understanding. Typical region-level benchmarks only evaluate the caption quality for single prompt using conventional language-based captioning metrics (You et al., 2023; Yuan et al., 2024; Zhang et al., 2024b; Guo et al., 2024; Rasheed et al., 2024), model-based similarities (Chen et al., 2025; Yuan et al., 2024), and LLM-Judged accuracies without the need for reference captions (Lian et al., 2025). GAR-Bench is to systematically evaluate the comprehension capabilities with multiple visual prompts. It contains a caption protocol to measure the correctness of descriptions for the relation between visual prompts, and a VQA protocol to evaluate both the basic understanding capability for specific regions, e.g., color and shape, and advanced compositional reasoning abilities for multiple regions.
@@ 3 Grasp Any Region
We start from the task formulation in Section 3.1. Subsequently, we introduce our model architecture and training data pipeline in Section 3.2 and Section 3.3, respectively. Finally, we introduce our benchmark designs in Section 3.4 to systematically evaluate region-level comprehension capabilities.
@@ 3.2 Model Architecture
Prompt Encoding and Integration. To integrate spatial guidance into the vision backbone, we introduce a lightweight prompt encoding mechanism similar to Lian et al. (2025) and Sun et al. (2024). The input binary mask, which specifies the region(s) of interest, is first processed by a simple convolutional block (LeCun et al., 1989) to produce a mask embedding. This zero-initialized (Zhang et al., 2023) mask embedding is then added to ViT’s (Dosovitskiy et al., 2021) patch embeddings.
RoI-aligned Feature Replay. To simultaneously provide sufficient local details and maintain necessary global context, we introduce the RoI-aligned feature replay technique. Specifically, our model processes the full, uncropped image (with the encoded mask prompt) with AnyRes (Liu et al., 2024), producing a global feature map that is rich in contextual information. Based on the input mask, we then derive a corresponding bounding box for the region of interest and employ RoI-Align (He et al., 2017) to gather the relevant feature vectors directly from the global feature map. Because the features are extracted from a feature map that was computed over the entire image, they are inherently context-aware, which elegantly avoids the pitfalls of local-only processing in Lian et al. (2025). At the same time, it provides the subsequent language model with a sufficiently detailed, high-resolution representation of the prompted region, enabling it to perform fine-grained understanding. This replay of context-rich features allows GAR to simultaneously “zoom in” on detail without “losing sight” of the bigger picture. Ablations of this design can be found in Table 9, where we demonstrate that this design is capable of both (1) providing sufficient local details and (2) preserving global contexts.
@@ 3.3 Training Data Pipeline
To enhance model capabilities from basic object recognition with single region to complex relational reasoning with multiple regions, we design a multi-stage process to generate a large-scale, high-quality dataset, as illustrated in Figure 4. Ablations of each round can be found in Table 11. Prompts for each stage can be found in Appendix G.
Round 1: Enhance Recognition Capability. Initially, we start from the Describe Anything-1.5M dataset (Lian et al., 2025). However, we observe deficiencies in its fine-grained recognition capability, limiting the quality of generated captions for more complex scenarios. To address this, we integrated images and masks provided by Sun et al. (2024), which is a subset of ImageNet-21K (Deng et al., 2009), an extremely fine-grained classification dataset and renowned for its detailed and extensive category labels. We employ the seed captioner to generate descriptions and then utilize an LLM to validate these generated captions against the ground-truth categories, resulting in a refined fine-grained dataset of 456K samples. We utilize both datasets to train a fine-grained captioner.
Round 2: Supporting Multiple Prompts. To further enable understanding multiple prompts, we incorporated the Panoptic Scene Graph (PSG) dataset (Yang et al., 2022), which is rich in relational information. We first query the fine-grained captioner to generate a detailed description for each region. Subsequently, we regard Qwen2.5-72B (Team, 2024) as the LLM-Merger, together with the original annotations provided by the PSG dataset (Yang et al., 2022), to generate: (1) 144K rich object descriptions that explicitly integrate relational context, (2) 144K question-answering pairs designed to probe the understanding of complex relationships, and (3) 126K multiple-choice questions. We construct a relation dataset with 414K samples in total during this stage.
@@ 4 Experiments
Owing to page limitations, we only present the key properties in this section. For implementation details, comparative baselines, and ablation studies, please refer to Appendix C.
@@
| Qwen2.5-VL-3B | 34.4 | 29.0 | 25.0 | 34.5 | 30.6 | 43.8 | 26.2 | 44.6 | ||||||||||||
| Qwen2.5-VL-7B | 41.7 | 39.1 | 40.6 | 44.8 | 27.8 | 59.4 | 36.1 | 40.6 | ||||||||||||
| Qwen2.5-VL-32B | 50.9 | 46.4 | 53.1 | 41.4 | 30.6 | 71.9 | 36.1 | 58.4 | ||||||||||||
| Qwen2.5-VL-72B | 52.8 | 46.4 | 50.0 | 65.5 | 33.3 | 68.8 | 44.3 | 57.4 | ||||||||||||
| InternVL3-2B | 35.1 | 30.4 | 21.9 | 48.3 | 38.9 | 48.4 | 26.2 | 38.6 | ||||||||||||
| InternVL3-8B | 38.9 | 36.2 | 37.5 | 58.6 | 41.7 | 51.6 | 27.9 | 33.6 | ||||||||||||
| InternVL3-38B | 46.5 | 39.1 | 40.6 | 51.7 | 55.6 | 60.9 | 36.1 | 47.5 | ||||||||||||
| InternVL3-78B | 50.5 | 44.9 | 54.7 | 58.6 | 61.1 | 53.1 | 47.5 | 45.5 | ||||||||||||
@@
| Qwen2.5-VL-3B | 22.5 | 9.3 | 34.6 | ||||||
| Qwen2.5-VL-7B | 32.4 | 12.4 | 50.5 | ||||||
| Qwen2.5-VL-32B | 36.8 | 17.5 | 54.3 | ||||||
| InternVL3-2B | 29.4 | 14.4 | 43.0 | ||||||
| InternVL3-8B | 33.8 | 11.3 | 54.2 | ||||||
| InternVL3-38B | 45.1 | 29.9 | 58.9 | ||||||
@@
| Shikra-7B | 22.2 | 2.7 | 41.8 |
| Ferret-7B | 22.4 | 6.4 | 38.4 |
| RegionGPT-7B | 27.2 | 13.0 | 41.4 |
| VP-SPHINX-13B | 22.5 | 11.7 | 33.2 |
| DAM-3B | 64.5‡ | 47.2‡ | 81.8‡ |
@@
Advanced comprehension requires precisely modeling complex relationships between multiple prompts. To evaluate this capability, we conducted a comprehensive comparison on our GAR-Bench-VQA. As demonstrated in Table 1, GAR-8B achieves an impressive overall score of 54.5, surpassing even the powerful, private, state-of-the-art non-thinking model, GPT-4o (OpenAI, 2024a). Furthermore, the efficiency and effectiveness of our approach are highlighted by GAR-1B. Despite its significantly smaller size, it scores 50.6 overall, outperforming large-scale public models like InternVL3-78B (Zhu et al., 2025). This advantage is particularly evident in fine-grained perception tasks, where GAR-1B and GAR-8B achieve “Texture” scores of 69.0 and 75.9, respectively.
@@
| LLaVA-OneVison-7B | 2.12 | 2.62 | 1.58 | 2.19 | 2.07 | 2.48 | 3.09 | 1.94 | 2.50 | 2.41 |
| Qwen2-VL-7B | 2.39 | 2.97 | 2.24 | 2.03 | 2.31 | 2.55 | 3.30 | 2.54 | 2.22 | 2.12 |
| InternVL2-26B | 2.84 | 3.55 | 2.99 | 2.57 | 2.25 | 3.20 | 4.08 | 3.35 | 3.08 | 2.28 |
@@
| Elysium-7B | 1.57 | 2.35 | 0.30 | 0.02 | 3.59 | – | – | – | – | – |
| Ferret-7B | 2.18 | 3.08 | 2.01 | 1.54 | 2.14 | 2.23 | 3.20 | 2.38 | 1.97 | 1.38 |
| Osprey-7B | 2.34 | 3.19 | 2.16 | 1.54 | 2.45 | 2.41 | 3.30 | 2.66 | 2.10 | 1.58 |
| Artemis-7B | – | – | – | – | – | 2.26 | 3.42 | 1.34 | 1.39 | 2.90 |
| DAM-8B | – | – | – | – | – | 3.34 | 4.45 | 3.30 | 3.03 | 2.58 |
@@
Detailed localized captioning requires generating detailed descriptions for given regions with multiple sentences. We benchmark our GAR models on a series of challenging datasets, and the results consistently demonstrate their state-of-the-art capabilities. As shown in Table 3, on our GAR-Bench-Cap, GAR-1B and GAR-8B achieve the highest overall scores of 57.5 and 62.2, respectively, even exceeding that of powerful private models like Gemini-2.5-Pro (DeepMind, 2025b). This superiority is further confirmed on the DLC-Bench (Lian et al., 2025) in Table 3, where GAR-1B and GAR-8B again outperform top models like DAM-3B using either LLaMA3.1 (Grattafiori et al., 2024) or GPT-4o (OpenAI, 2024a) as the judge. The zero-shot performance of our models on Ferret-Bench (You et al., 2023) and MDVP-Bench (Lin et al., 2025a), detailed in Table 5, is particularly noteworthy. On both benchmarks, our GAR emerges as the top-performing model across every single category. Specifically on MDVP-Bench, our models show a commanding lead, with GAR-8B achieving a score of 178.6 on natural images, a result that is substantially higher than any competitor. Collectively, these comprehensive evaluations across multiple benchmarks unequivocally establish GAR as the new state-of-the-art for producing rich, accurate, and detailed localized captions.
Open-class category-level image recognition requires the model to recognize the category of the object and part entities. We evaluate this capability in Table 5. Our GAR-8B demonstrates a significant leap in performance, establishing a new state-of-the-art. It consistently outperforms all prior methods across every metric, achieving top scores of 93.6 semantic similarity and 88.7 semantic IoU on LVIS (Gupta et al., 2019), and 95.5 semantic similarity and 91.8 semantic IoU on PACO (Ramanathan et al., 2023). This indicates its superior ability in both semantic understanding and precise localization. These results demonstrate the effectiveness of GAR for complex recognition tasks, showcasing its robust performance in identifying a diverse range of object categories.
Extension to videos is straightforward. Similar to Lian et al. (2025), we simply extend our GAR models to videos and evaluate them on VideoRefer-Bench (Yuan et al., 2025b) and VideoRefer-Bench (Yuan et al., 2025b) in Table 6 and Table 7, respectively. We uniformly sample 16 frames to represent a video. Our GAR-8B surpasses DAM-8B (Lian et al., 2025) under the zero-shot setting. More importantly, as demonstrated in Table 7, our our zero-shot GAR-8B even outperforms in-domain VideoRefer-7B, demonstrating its strong comprehension capabilities can be easily transferred to videos. However, as our models are actually trained with images, they get reasonably low scores on temporally related tasks, e.g., temporal description (TD) in Table 6 and future predictions in Table 7.
@@
| InternVL2-26B | 65.0 | 58.5 | 63.5 | 53.4 | 88.0 | 78.9 | |||||||||||||||||
| Qwen2-VL-7B | 66.0 | 62.0 | 69.6 | 54.9 | 87.3 | 74.6 | |||||||||||||||||
| LLaVA-OneVision-7B | 67.4 | 58.7 | 62.9 | 64.7 | 87.4 | 76.3 | |||||||||||||||||
| GPT-4o | 71.3 | 62.3 | 74.5 | 66.0 | 88.0 | 73.7 | |||||||||||||||||